OpenStudy (youngster):

Try to Create a good model to describe this growth: 1991-----27 1992-----36 1993-----45 1994-----181 1995-----306 1996-----1093 1997-----2459

OpenStudy (youngster):

1991-----27 1992-----36 1993-----45 1994-----181 1995-----306 1996-----1093 1997-----2459

OpenStudy (anonymous):

by first guess would be to take the ration of \(\frac{36}{27}=\frac{4}{3}\) and see if it looks like every year the populations grows by factor of \(\frac{4}{3}\)

OpenStudy (youngster):

I know im suppose to somehow use the stats functions on my graph calc

OpenStudy (anonymous):

oh, in any case i was wrong that was just for the first year, growth changes after that

OpenStudy (youngster):

but we're supposed to do something to make it linear, then work from that

OpenStudy (anonymous):

this sure as hell is not linear

OpenStudy (youngster):

we're supposed to make the x axis log or somthing...